# Twists of Superconformal Algebras

### v2

@@ 4.3. Complex Group Orbits
This map is locally a diffeomorphism. Indeed, working in a neighborhood of matrices $(A,B)$ where $\det(A)$ lies in a small disk in $\mathbb{C}^{\times}$, and choose a local fourth root in this disk, we can define a local inverse sending $(A,B)$ to $(A/\det(A),B/\det(B),\det(A)^{1/4})$.
@@
When both terms $\mathcal{Q}_{\pm}$ have positive rank, the orbit of the twist $(\mathcal{Q}_{+},\mathcal{Q}_{-})$ is an isolated point. When one term has rank 0 and the other has rank 4 (i.e., $(r_{+},r_{-})=(4,0)$ or $(0,4)$), then the orbit sits in a one-parameter family labeled by the determinant of the full rank term (in the $(4,0)$ case, $\det(\mathcal{Q}_{+})$).
@@ 4.4. Closed and Exact Transformations
    Rank $(0,0)$: The zero supercharge always squares to zero. If $k<2$ it is the only point in $\mathrm{Nilp}_{(4|k)}(\mathbb{C})$. The kernel of $[\mathcal{Q},-]$ is $\mathfrak{so}(6,\mathbb{C})\oplus\mathfrak{sl}(k,\mathbb{C})\oplus\mathbb{C}$ (the entire even component), and the image of $[\mathcal{Q},-]$ is just $\{0\}$.
@@
    $\mathfrak{b}_{r}$ =$=\Bigg{\{}\left(\left(\begin{array}[]{c|c}\alpha&\beta\\ \hline\cr 0&0\end{array}\right),\left(\begin{array}[]{c|c}\alpha&0\\ \hline\cr\gamma&0\end{array}\right),\lambda\right)\in\mathfrak{sl}(4,\mathbb{C})\oplus\mathfrak{sl}(k,\mathbb{C})\oplus\mathbb{C}\colon$
@@ 1 $k^{2}-k+11+\delta_{kr}-\delta_{k4}$ $k+3$ $k^{2}-2k+8+\delta_{kr}-\delta_{k4}$
    2 $k^{2}-2k+10+\delta_{kr}-\delta_{k4}$ $42k+$ 4 $k^{2}-4k+6+\delta_{kr}-\delta_{k4}$
    3 $k^{2}-3k+11+\delta_{kr}-\delta_{k4}$ $33k+$ 3 $k^{2}-6k+8+\delta_{kr}-\delta_{k4}$
@@ 4 $k^{2}-4k+14+\delta_{kr}-\delta_{k4}$ $4k$ $k^{2}-8k+14+\delta_{kr}-\delta_{k4}$
    Table 2. For chiral supercharges $\mathcal{Q}$ of rank $r$, we give the dimension of the kernel and image of $[\mathcal{Q},-]$ as well as the dimension of the quotient.
    We note from this computation that the inclusion $\mathfrak{z}_{\mathcal{Q}}\subseteq\mathfrak{b}_{\mathcal{Q}}$ is always proper, at least for $k\neq 4$. In the $k=\mathcal{N}=4$ case the computation remains the same, but $\dim\mathfrak{z}_{\mathcal{Q}}$ is decreased by one. The inclusion $\mathfrak{z}_{\mathcal{Q}}\subseteq\mathfrak{b}_{\mathcal{Q}}$ is still proper, with the exception of the maximal rank $r=4$ case where $\mathfrak{b}_{\mathcal{Q}}=\mathfrak{z}_{\mathcal{Q}}$ and they span the diagonal copy of $\mathfrak{sl}(4,\mathbb{C})\subseteq\mathfrak{sl}(4,\mathbb{C})\oplus\mathfrak{sl}(4,\mathbb{C})$.
@@
    Let $\mathcal{Q}=Q_{+}\otimes w_{+}+Q_{-}\otimes w_{-}$ be a rank $(1,1)$ square zero supercharge. The image $\mathfrak{b}_{\mathcal{Q}}\subseteq\mathfrak{sl}(4,\mathbb{C})\times\mathfrak{sl}(k,\mathbb{C})\times\mathbb{C}$ of the linear map $[\mathcal{Q},-]$ is isomorphic to
@@
    One can check both propositions using the same setup. If $\{Q_{1},Q_{2},Q_{3},Q_{4}\}$ and $\{w_{1},\ldots,w_{k}\}$ are bases for $S_{+}$ and $W$, respectively, without loss of generality we may let $\mathcal{Q}=Q_{1}\otimes w_{1}+Q_{2}^{*}\otimes w_{2}^{*}$. In this basis $\mathcal{Q}$ is represented by a $k\times 4$ matrix and a $4\times k$ matrix with ones in the $(1,1)$ and $(2,2)$ positions respectively, and zeroes in all other positions. It is then an elementary computation to work out the image and kernel of the operator $[\mathcal{Q},-]$.
@@
Of course the classification does not need to end here: there are conformal twists of higher rank, including twists of rank $(2,2)$ generalizing the family of twists in the $\mathcal{N}=4$ super Poincaré algebra studied by Kapustin and Witten [KapustinWitten]. We leave a detailed analysis of all cases for future work.
@@ 5. Chiral Twists in 4d Superconformal Algebras: Some Case by Case Analysis
In this section we discuss the orbit structure of superconformal twists under the action of real superconformal groups in various signatures. We stick to the chiral twists as the situation is already complicated there, and we eventually specialize to the situation that Beem et al [Beem_2015] showed to be strikingly rich.
@@ 5.1. Real Orbits of Chiral Twists: General Features
where on the far right, we use the natural action of $B\in\mathrm{GL}(2,\mathbb{C})$ on the projective line. The action of the factor $A\in\mathrm{SL}(4,\mathbb{C})$ preserves the fiber $\pi^{-1}(L)$ for each line $L\subset\mathbb{C}^{2}$, since $A\mathcal{Q}B$ has the same kernel as $\mathcal{Q}$. (The image of $\mathcal{Q}$ changes, however, under the action of $\mathrm{SConf}(4|2,\mathbb{C})_{0}$.)
@@
In the case where $k=2,r=1$ discussed in Example 5.1 above we realized $\mathbb{O}(4|2)_{(1,0)}$ as a fiber bundle over $\mathbb{CP}^{1}$ with fiber isomorphic to the space of rank one $2\times 4$ matrices with fixed kernel. We can identify this fiber with $\mathbb{C}^{4}\backslash\{0\}$ with its natural action of $\mathrm{SL}(4,\mathbb{C})$. We will discuss the real forms of $\mathrm{SConf}(4|2,\mathbb{C})_{0}$ in general in a moment, but we may summarize here what happens in each of the three possible signatures for $\mathbb{R}^{p,4-p}$.
@@
    In signature $(4,0)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SL}(2,\mathbb{H})$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SU}(2)$. The action of $\mathrm{SU}(2)$ on $\mathbb{CP}^{1}$ is transitive, as is the action of $\mathrm{SL}(2,\mathbb{H})$ on $\mathbb{C}^{4}\backslash\{0\}\cong\mathbb{H}^{2}\backslash\{0\}$. In this case the real orbit space is a single point. (That is, there is a single orbit; all nilpotent supercharges are equivalent.)
@@
    In signature $(3,1)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SU}(2,2)$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SU}(2)$. The action of $\mathrm{SU}(2)$ on $\mathbb{CP}^{1}$ is transitive, but the action of $\mathrm{SL}(2,\mathbb{H})$ on $\mathbb{C}^{4}\backslash\{0\}$ is more interesting: there is a continuous family of orbits parameterized by $\mathbb{R}^{\times}$ (corresponding to non-zero values of the split signature Hermitian form on $\mathbb{C}^{2,2}$), but there is also a family of orbits with a finite non-Hausdorff locus (corresponding to the null locus for the Hermitian form).
    That is, the quotient space by the $\mathrm{SL}(2,\mathbb{H})$-action is non-Hausdorff.
@@
    In signature $(2,2)$ we have $G_{\mathbb{R}}\cap\mathrm{SL}(4,\mathbb{C})\cong\mathrm{SL}(4,\mathbb{R})$, and $G_{\mathbb{R}}\cap\mathrm{SL}(2,\mathbb{C})\cong\mathrm{SL}(2,\mathbb{R})$. The action of $\mathrm{SL}(2,\mathbb{R})$ on $\mathbb{CP}^{1}$ has three orbits: two open orbits (the two open hemispheres) and one closed orbit (the equator), so the quotient space has three points with a connected non-Hausdorff topology in which a single point is dense. The action of $\mathrm{SL}(4,\mathbb{R})$ on $\mathbb{C}^{4}\backslash\{0\}$ is again interesting, because the quotient space parametrizing orbits is a connected non-Hausdorff space containing an open subspace isomorphic to $\mathbb{R}^{\times}$.
@@
Theorem 5.5 ([Wolf, Theorem 2.6], [FHW, Theorem 3.2.1]).
@@
In our case of interest, $n=4$ and $m=4-r$. The theorem implies that there are finitely many orbits in the bottom right corner:
@@
The number $c$ appearing in the upper bound for the number of orbits $|D^{(r)}_{*,*}|$ is computed for general real forms of $\mathrm{SL}(n,\mathbb{C})$ by Sugiura [Sugiura]. In particular we find $c=1$ in Euclidean signature and $c=3$ in split signature. The remaining Lorentzian signature case is a hermitian symmetric space so we obtain the exact number of orbits as being 6.
@@ 5.2.1. Signature $(4,0)$
Given a subspace $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, representing a point $[V]\in\mathrm{Gr}(2,4)$, we obtain an $\mathbb{H}$-linear subspace $V_{\mathbb{H}}\subset\mathbb{H}^{2}$ by taking the $\mathbb{H}$-span of $V$, i.e.,
@@ 5.2.3. Signature $(2,2)$
Let $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, and hence represent a point $[V]\in\mathrm{Gr}(2,4)$. We will study its intersection with $\mathbb{R}^{4}$ (the “real intersection”).
@@
$\mathrm{GL}(2,\mathbb{C})/(\mathrm{SL}(2,\mathbb{R})\times\mathbb{R}^{\times})\cong\mathcal{(}\mathbb{R}^{2}\times(S^{1})^{2},$
@@ 5.3. The Case of 4d $\mathcal{N}=2$ Chiral Supercharges
In this section we will analyze the specific example considered by Beem et al [Beem_2015] in their derivation of vertex algebras from 4d $\mathcal{N}=2$ superconformal field theory. Beem et al are concerned with a chiral supercharge of rank $(2,0)$ that they dub the twisted Schur supercharge. The analysis above tells us that for this example
@@
For example, in this basis, the supercharge $\mathcal{Q}_{+1}$ is a Donaldson–Witten supercharge [Witten_1988]. Note that it lies in the first summand $S_{+}\otimes W_{2}$, and hence in an ordinary 4d $\mathcal{N}=2$ super Poincaré algebra. Thus it acts on supersymmetric theories, not just superconformal theories, and famously twists $\mathcal{N}=2$ super Yang–Mills theory into a 4d topological field theory.
@@
Beem et al [Beem_2015] refer to any rank $(2,0)$ elements the do not lie in either summand as Schur supercharges. The name comes from their connection to the Schur limit of the superconformal index of 4d $\mathcal{N}=2$ superconformal theories. (See the introduction and Section 4.4 of loc. cit. for explanation of this context.) Following their terminology, a “Schur operator” in a superconformal theory is a local operator representing a non-trivial cohomology class for a fixed Schur supercharge.
Beem al discovered that these Schur operators yield vertex algebras, providing a beautiful and powerful relationship between 4d and 2d field theory. We return to this aspect in Section 8, where we explain how conformal twists modify observables of superconformal field theories. In particular, we discuss in Section 8.4 some powerful results of Saberi and Williams on this topic that provide many examples of how factorization techniques and twisting combine to yield vertex algebras, which are complementary to the structural work we do in Section 8.
@@ 5.3.1. The real forms of the cohomology of a Schur supercharge
$\pi(\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}})$ =$=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{p}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
$\pi(\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}})$ =$=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{k}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
@@
Remark 5.15.
One can compare this discussion to the explicit description of the closed and exact transformations appearing in [Argyres:2022npi, Section 2.3]. The generators of $\mathfrak{z}_{\mathcal{Q}}^{\mathbb{C}}$ are denoted there by $\mathfrak{sl}_{2}\oplus\widehat{\mathfrak{sl}_{2}}\oplus\mathcal{Z}\oplus\mathfrak{m}_{+}$, and the factor $\mathfrak{m}_{+}$ is orthogonal to the fixed point locus of the involution $\theta_{\mathrm{Lor}}$.

### v3

@@ 4.3. Complex Group Orbits
This map is locally a diffeomorphism. Indeed, taking in a neighborhood of matrices $(A,B)$ where $\det(A)$ lies in a small disk in $\mathbb{C}^{\times}$ and choosing a local fourth root in this disk, we can define a local inverse sending $(A,B)$ to $(A/\det(A),B/\det(B),\det(A)^{1/4})$.
@@
When both terms $\mathcal{Q}_{\pm}$ have positive rank, the orbit of the twist $(\mathcal{Q}_{+},\mathcal{Q}_{-})$ is an isolated point. When one term has rank 0 and the other has rank 4 (i.e., $(r_{+},r_{-})=(4,0)$ or $(0,4)$), then the orbit sits in a one-parameter family labeled by the determinant of the full rank term (e.g., by $\det(\mathcal{Q}_{+})$ in the $(4,0)$ case).
@@ 4.4. Closed and Exact Transformations
    Rank $(0,0)$: The zero supercharge always squares to zero. The kernel of $\mathcal{Q},-]$ is $\mathfrak{so}(6,\mathbb{C})\oplus\mathfrak{sl}(k,\mathbb{C})\oplus\mathbb{C}$ (the entire even component), and the image of $\mathcal{Q},-]$ is just $\{0\}$.
@@
    $\mathfrak{b}_{r}$ $=\Bigg\{\left(\left(\begin{array}[]{c|c}\alpha&\beta\\ \hline\cr 0&0\end{array}\right),\left(\begin{array}[]{c|c}\alpha&0\\ \hline\cr\gamma&0\end{array}\right),\lambda\right)\in\mathfrak{sl}(4,\mathbb{C})\oplus\mathfrak{sl}(k,\mathbb{C})\oplus\mathbb{C}\colon$
@@ 1 $k^{2}-k+11+\delta_{kr}-\delta_{k4}$ $k+3$ $k^{2}-2k+8+\delta_{kr}-\delta_{k4}$
    2 $k^{2}-2k+10+\delta_{kr}-\delta_{k4}$ $2k+4$ $k^{2}-4k+6+\delta_{kr}-\delta_{k4}$
    3 $k^{2}-3k+11+\delta_{kr}-\delta_{k4}$ $3k+3$ $k^{2}-6k+8+\delta_{kr}-\delta_{k4}$
@@ 4 $k^{2}-4k+14+\delta_{kr}-\delta_{k4}$ $4k$ $k^{2}-8k+14+\delta_{kr}-\delta_{k4}$
    Table 2. For chiral supercharges $\mathcal{Q}$ of rank $r$, we give the dimension of the kernel and image of $\mathcal{Q},-]$ as well as the dimension of the quotient.
    We note from this computation that the inclusion $\mathfrak{b}_{\mathcal{Q}}\subseteq\mathfrak{z}_{\mathcal{Q}}$ is always proper, at least for $k\neq 4$. In the $k=\mathcal{N}=4$ case the computation remains the same, but $\dim\mathfrak{z}_{\mathcal{Q}}$ is decreased by one. The inclusion $\mathfrak{b}_{\mathcal{Q}}\subseteq\mathfrak{z}_{\mathcal{Q}}$ is still proper, with the exception of the maximal rank $r=4$ case where $\mathfrak{b}_{\mathcal{Q}}=\mathfrak{z}_{\mathcal{Q}}$ and they span the diagonal copy of $\mathfrak{sl}(4,\mathbb{C})\subseteq\mathfrak{sl}(4,\mathbb{C})\oplus\mathfrak{sl}(4,\mathbb{C})$.
@@
    Let $\mathcal{Q}=Q_{+}\otimes w_{+}+Q_{-}\otimes w_{-}$ be a rank $(1,1)$ square zero supercharge. The image $\mathfrak{b}_{\mathcal{Q}}\subseteq\mathfrak{sl}(4,\mathbb{C})\times\mathfrak{sl}(k,\mathbb{C})\times\mathbb{C}$ of the linear map $\mathcal{Q},-]$ is isomorphic to
@@
    One can check both propositions using the same setup. If $\{Q_{1},Q_{2},Q_{3},Q_{4}\}$ and $\{w_{1},\ldots,w_{k}\}$ are bases for $S_{+}$ and $W$, respectively, without loss of generality we may let $\mathcal{Q}=Q_{1}\otimes w_{1}+Q_{2}^{*}\otimes w_{2}^{*}$. In this basis $\mathcal{Q}$ is represented by a $k\times 4$ matrix and a $4\times k$ matrix with ones in the $(1,1)$ and $(2,2)$ positions respectively, and zeroes in all other positions. It is then an elementary computation to work out the image and kernel of the operator $\mathcal{Q},-]$.
@@
Of course the classification does not need to end here: there are conformal twists of higher rank, including twists of rank $(2,2)$ generalizing the family of twists in the $\mathcal{N}=4$ super Poincaré algebra studied by Kapustin and Witten [32]. We leave a detailed analysis of all cases for future work.
@@ 5. Chiral Twists in 4d Superconformal Algebras: Some Case by Case Analysis
In this section we discuss the orbit structure of superconformal twists under the action of real superconformal groups in various signatures. We stick to the chiral twists as the situation is already complicated there, and we eventually specialize to the situation that Beem et al [6] showed to be strikingly rich.
@@ 5.1. Real Orbits of Chiral Twists: General Features
where on the far right, we use the natural action of $B\in\mathrm{GL}(2,\mathbb{C})$ on the projective line. The action of the factor $A\in\mathrm{SL}(4,\mathbb{C})$ preserves the fiber $\pi^{-1}(L)$ for each line $L\subset\mathbb{C}^{2}$, since $A\mathcal{Q}B$ has the same kernel as $\mathcal{Q}B$. (The image of $\mathcal{Q}$ changes, however, under the action of $\mathrm{SConf}(4|2,\mathbb{C})_{0}$.)
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Theorem 5.5 ([46, Theorem 2.6], [24, Theorem 3.2.1]).
@@
In our case of interest, $n=4$ and m=4rm=4-r. The theorem implies that there are finitely many orbits in the bottom right corner:
@@
The number $c$ appearing in the upper bound for the number of orbits $|D^{(r)}_{*,*}|$ is computed for general real forms of $\mathrm{SL}(n,\mathbb{C})$ by Sugiura [44]. In particular we find $c=1$ in Euclidean signature and $c=3$ in split signature. The remaining Lorentzian signature case is a hermitian symmetric space so we obtain the exact number of orbits as being 6.
@@ 5.2.1. Signature $(4,0)$
Given a subspace $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, representing a point $V]\in\mathrm{Gr}(2,4)$, we obtain an $\mathbb{H}$-linear subspace $V_{\mathbb{H}}\subset\mathbb{H}^{2}$ by taking the $\mathbb{H}$-span of $V$, i.e.,
@@ 5.2.3. Signature $(2,2)$
Let $V\subset\mathbb{C}^{4}$ with $\dim_{\mathbb{C}}(V)=2$, and hence represent a point $V]\in\mathrm{Gr}(2,4)$. We will study its intersection with $\mathbb{R}^{4}$ (the “real intersection”).
@@
$\mathrm{GL}(2,\mathbb{C})/(\mathrm{SL}(2,\mathbb{R})\times\mathbb{R}^{\times})\cong\mathcal{\mathbb{R}}^{2}\times(S^{1})^{2},$
@@ 5.3. The Case of 4d $\mathcal{N}=2$ Chiral Supercharges
In this section we will analyze the specific example considered by Beem et al [6] in their derivation of vertex algebras from 4d $\mathcal{N}=2$ superconformal field theory. Beem et al are concerned with a chiral supercharge of rank $(2,0)$ that we will call the Schur supercharge. The analysis above tells us that for this example
@@
For example, in this basis, the supercharge $\mathcal{Q}_{+1}+\mathcal{Q}_{-2}$ is a Donaldson–Witten supercharge [45]. Note that it lies in the first summand $S_{+}\otimes W_{2}$, and hence in an ordinary 4d $\mathcal{N}=2$ super Poincaré algebra. Thus it acts on supersymmetric theories, not just superconformal theories, and famously twists $\mathcal{N}=2$ super Yang–Mills theory into a 4d topological field theory.
@@
Exploiting some features of the “Schur limit” of a 4d $\mathcal{N}=2$ superconformal theory [25], Beem et al [6] discovered how to obtain vertex algebras from 4d $\mathcal{N}=2$ superconformal theories, providing a beautiful and powerful relationship between 4d and 2d field theory. We return to this aspect in Section 8, where we explain how conformal twists modify observables of superconformal field theories. In particular, we discuss in Section 8.4 some powerful results of Saberi and Williams on this topic that provide many examples of how factorization techniques and twisting combine to yield vertex algebras, which are complementary to the structural work we do in Section 8.
@@ 5.3.1. The real forms of the cohomology of a Schur supercharge
$\pi(\mathfrak{z}_{\mathbb{Q}}^{\mathbb{C}})$ $=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{p}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$
$\pi(\mathfrak{b}_{\mathbb{Q}}^{\mathbb{C}})$ $=\left\{\left(\begin{array}[]{c|c}A&B\\ \hline\cr C&D\end{array}\right)\colon A,B,C,D\in\mathfrak{k}\text{ and }\mathrm{Tr}(A)+\mathrm{Tr}(D)=0\right\},$